# VideoScore: Building Automatic Metrics to Simulate Fine-grained Human Feedback for Video Generation

### v2

@@ Abstract
The recent years have witnessed great advances in video generation. However, the development of automatic video metrics is lagging significantly behind. None of the existing metric is able to provide reliable scores over generated videos. The main barrier is the lack of large-scale human-annotated dataset. In this paper, we release VideoFeedback, the first large-scale dataset containing human-provided multi-aspect score over 37.6K synthesized videos from 11 existing video generative models. We train VideoScore (initialized from Mantis) based on VideoFeedback to enable automatic video quality assessment. Experiments show that the Spearman correlation between VideoScore and humans can reach 77.1 on VideoFeedback-test, beating the prior best metrics by about 50 points. Further result on other held-out EvalCrafter, GenAI-Bench, and VBench show that VideoScore has consistently much higher correlation with human judges than other metrics. Due to these results, we believe VideoScore can serve as a great proxy for human raters to (1) rate different video models to track progress (2) simulate fine-grained human feedback in Reinforcement Learning with Human Feedback (RLHF) to improve current video generation models.
@@
1,2Xuan He1Dongfu Jiang1,3Ge Zhang  1Max Ku 1Achint Soni1Sherman Siu1Haonan Chen1Abhranil Chandra1Ziyan Jiang 1Aaran Arulraj4Kai Wang1Quy Duc Do1Yuansheng Ni2Bohan Lyu 1Yaswanth Narsupalli1Rongqi Fan1Zhiheng Lyu5Yuchen Lin1Wenhu Chen 1University of Waterloo, 2Tsinghua University, 3Stardust.AI, 4University of Toronto, 5AI2 Equal Contribution {x36he, dongfu.jiang, wenhuchen}@uwaterloo.ca
@@ 2.2 Video Quality Assessment
As the current progress of Text-to-Video generative models leaves it uncertain how close we are to reaching the objective, researchers have worked on evaluation methods to benchmark the generative models. Common methods involve the use of FVD Unterthiner et al. (2018) and CLIP Radford et al. (2021a) to evaluate the quality of frames and the text-frames alignment respectively. However, other aspects like subject consistency, temporal consistency, factualness cannot be captured by these metrics. Recent works like VBench Huang et al. (2023) proposes to use different DINO Caron et al. (2021), optical flow Horn and Schunck (1981) to reflect these aspects. However, the correlation with human judgment is relatively low. For example, most models have subject/background consistency scores over 97% in VBench, which is a massive overestimation of the current T2V models’ true capability. Another work EvalCrafter Liu et al. (2023b) instead resorts to human raters to perform comprehensive evaluation.
@@ 2.3 RLHF in image/video generation
In the recent years, reinforcement learning from human feedback (RLHF) has emerged as a significant approach in enhancing the performance of image/video generative models. Numerous studies have focused on training reward models with large datasets of image-text pairs, such as HPSv2 Wu et al. (2023b), ImageReward Xu et al. (2023), RichHF-18K Liang et al. (2023), or video-text pairs like T2V-Score Wu et al. (2024), while some research efforts have combined existing reward models, such as T2V-Turbo Li et al. (2024b), to simulate human feedback. Utilizing these reward models or feedback simulators, diverse methods have been proposed to align the output of visual generative models with human preferences, including RL-based methods (Fan et al. (2024), Zhang et al. (2024)) and reward fine-tuning methods (Clark et al. (2023), Li et al. (2024a), Yuan et al. (2023)). Additionally, some works adopts data distillation to fine-tune diffusions models on high-quality data, while others like Diffusion-DPO Wallace et al. (2023), extend Direct Preference Optimization (DPO) to train diffusion models based on preference data. Our VideoScore aims to approximate human feedback, which is expected to be beneficial in enhancing video generative models with RLHF methods like PPO or DPO.
@@ 4.5 Ablation Study
To investigate the effects of changing the base model, we have trained different variants with VideoLLaVA-7B and Idefics2-8B as the base models. Since VideoFeedback-test, EvalCrafter, and VBench both have multiple aspects in the benchmarks, we take the average score across these aspects and report the general performance in  Table 7. The results show that the Video-LLaVA-based version gets the worst performance on the four benchmarks, even if it is specifically designed for video understanding. The Idefics2-8B-based version has marginal improvements compared to the VideoLLaVA. After changing to Mantis-Idefics2-8B, the scores on the four benchmarks keep improving from $47.5$ to $55.6$ on average. When the scoring type is regression, the mantis-based version is still better than the Idefics2-based version by $12.1$ points. Therefore, we select the Mantis-based version as the final choice.
@@ Appendix F Annotation Details
Visual Quality
Expected Case:(1) The video looks clear and normal on its appearance.
(2) The features like Brightness, Contrast, Color, etc, are appropriate and stable.
Error point:(a) local obvious unclear or blurry,
(b) too low resolution,
(c) some speckles or black patches,
(d) appearance of video is skewed and distorted,
(e) unstable optical property, such as brightness, contrast, saturation, exposure etc,
(f) flickering color of main objects and background
Note:Some videos have watermark, we can ignore that.
Temporal Consistency
Expected Case:
(1) The main objects, main characters and overall appearance are consistent
across the video.
(2) The appearance of video as well as the movements of humans and objects
are smooth and natural.
Error points:
(a) The person or object suddenly disappears or appears
(b) The type or class of objects has obvious changes
(c) There is an obvious switch in the screen shot
(d) the appearance of video or movements in it is laggy and un-smooth,
(e) local deformation or dislocation of human or objects due to the motion
(for large scale deformation, the video should also be rated as bad in
"1. visual quality"),
Note:
For a video almost static or with small dynamic degree, as long as it does not
have error points, then
it should be scored as good.
Dynamic Degree
Expected Case:
(1) The video is obviously not static, the people or objects or the video screen
is dynamic.
(2) The video can be easily distinguished from a static image.
Note:
You are supposed to focus on only dynamic degree, regardless of the visual
quality and video content
Text-to-Video Alignment
Expected Case:
The characters, objects, motions, events etc. that are mentioned in text input
prompts all exist reasonably.
Error points:
(a) The people and objects in prompt do not appear in video
(b) The actions and events in prompt do not appear in video
(c) The number, size, shape, color, state, movement and other attributes of
the objects in the video do not match the prompt
(d) Text mentioned in prompt is not displayed correctly in the video,
such as "a placard saying ’No Smoking’" but "No Smoking" is not spelled
correctly in the video
(e) The video format (such as width, height, screen ratio, duration) does not
match the format in prompt.
Factual Consistency
Expected Case:
(1) Overall appreance and motion are consistent with our common-sense,
physical principles, moral standards, etc.
Error points:
(a) static ones: Content in video goes against common sense in life, such as
lighting a torch in the water, standing in the rain but not getting wet, etc.
(b) static ones: The size, color, shape and other basic properties of objects
violate scientific principles
(c) dynamic ones: The overall movement of people or objects violates
common-sense and laws of physics, such as spontaneous upward movement
against gravity, abnormal water flow, etc.
(d) dynamic ones: Partial movements of people or objects violate common-sense
and laws of physics, such as the movement of hands or legs is anti-joint, etc.
Notes:
Relation with ’5. text-to-video alignment’:
Some text prompts express fictional and unrealistic content, for example,
"a dog plays the guitar in the sky" or "an astronaut rides a horse in space".
In this case, regardless of the veracity of the text prompt, you should
only consider whether the other content in the video makes sense.

### v3

@@ Abstract
The recent years have witnessed great advances in video generation. However, the development of automatic video metrics is lagging significantly behind. None of the existing metrics are able to provide reliable scores over generated videos. The main barrier is the lack of large-scale human-annotated datasets. In this paper, we release VideoFeedback, the first large-scale dataset containing human-provided multi-aspect score over 37.6K synthesized videos from 11 existing video generative models. We train VideoScore (initialized from Mantis) based on VideoFeedback to enable automatic video quality assessment. Experiments show that the Spearman correlation between VideoScore and humans can reach 77.1 on VideoFeedback-test, beating the prior best metrics by about 50 points. Further results on other held-out EvalCrafter, GenAI-Bench, and VBench show that VideoScore has consistently much higher correlation with human judges than other metrics. Due to these results, we believe VideoScore can serve as a great proxy for human raters to (1) rate different video models to track progress (2) simulate fine-grained human feedback in Reinforcement Learning with Human Feedback (RLHF) to improve current video generation models.
@@
1,2Xuan He1Dongfu Jiang1,3Ge Zhang  1Max Ku 1Achint Soni   1Sherman Siu   1Haonan Chen   1Abhranil Chandra   1Ziyan Jiang 1Aaran Arulraj   4Kai Wang   1Quy Duc Do   1Yuansheng Ni   2Bohan Lyu 1Yaswanth Narsupalli  1Rongqi Fan  1Zhiheng Lyu  5Bill Yuchen Lin  1Wenhu Chen 1University of Waterloo, 2Tsinghua University, 3Stardust.AI, 4University of Toronto, 5AI2 Equal Contribution {x36he, dongfu.jiang, wenhuchen}@uwaterloo.ca
@@ 2.2 Video Quality Assessment
As the current progress of Text-to-Video generative models leaves it uncertain how close we are to reaching the objective, researchers have worked on evaluation methods to benchmark the generative models. Common methods involve the use of FVD Unterthiner et al. (2018) and CLIP Radford et al. (2021a) to evaluate the quality of frames and the text-frames alignment respectively. However, other aspects like subject consistency, temporal consistency, factual consistency cannot be captured by these metrics. Recent works like VBench Huang et al. (2023) proposes to use different DINO Caron et al. (2021), optical flow Horn and Schunck (1981) to reflect these aspects. However, the correlation with human judgment is relatively low. For example, most models have subject/background consistency scores over 97% in VBench, which is a massive overestimation of the current T2V models’ true capability. Another work EvalCrafter Liu et al. (2023b) instead resorts to human raters to perform comprehensive evaluation.
@@ 2.3 RLHF in image/video generation
In recent years, reinforcement learning from human feedback (RLHF) has emerged as a significant approach to enhancing the performance of image/video generative models. Numerous studies have focused on training reward models with large datasets of image-text pairs, such as HPSv2 Wu et al. (2023b), ImageReward Xu et al. (2023), RichHF-18K Liang et al. (2023), or video-text pairs like T2V-Score Wu et al. (2024). Moreover, some research efforts have combined existing reward models to simulate human feedback, such as T2V-Turbo Li et al. (2024c), while some recent works (Wang et al. (2024b), Wang et al. (2024a)) proposed multi-objective reward model with regression head to provide human preferences. Utilizing these reward models or feedback simulators, diverse methods have been proposed to align the output of visual generative models with human preferences, including RL-based methods (Fan et al. (2024), Zhang et al. (2024)) and reward fine-tuning methods (Clark et al. (2023), Li et al. (2024b), Yuan et al. (2023)). Additionally, some works adopt data distillation to fine-tune diffusion models on high-quality data, while others like Diffusion-DPO Wallace et al. (2023), extend Direct Preference Optimization (DPO) to train diffusion models based on preference data. Our VideoScore aims to approximate human feedback, which is expected to be beneficial in enhancing video generative models with RLHF methods like PPO or DPO.
@@ 4.4 Evaluation Results
4.5 Best-of-K Sampling with VideoScore
While best-of-K sampling has proven effective in boosting the performance of LLMs Li et al. (2023), it’s still unknown whether it also works for video generation tasks. To investigate this problem and also better show the effectiveness of VideoScore, we conduct a comparison of different T2V models with and without employing best-of-k sampling with VideoScore on EvalCrafter.
We set $k=5$ and generated videos using 700 prompts from EvalCrafter. For each prompt, the video with the highest average VideoScore across five dimensions was selected. We then evaluated both the "best videos" and randomly chosen ones using EvalCrafter’s metrics, averaging the results over 700 videos to obtain the model’s final score. As shown in  Table 8, compared to the random sample, most scores on the EvalCrafter benchmark have increased after the best-of-5 process.
T2V Model Dimensions of EvalCrafter
Average Visual Temporal Motion T2V Align
random best random best random best random best random best
AnimateDiff 62.16 61.87 60.79 60.89 60.69 60.94 54.35 54.83 72.79 70.83
HotShot-XL 53.69 61.22 57.56 60.39 58.85 60.94 51.52 54.05 46.83 69.52
LaVie-base 57.88 59.90 57.79 58.52 54.21 57.70 52.51 53.53 66.99 69.86
VideoCrafter2 58.31 58.98 58.44 59.70 59.18 60.79 54.39 54.65 61.23 60.77
VideoCrafter1 54.30 57.28 52.40 53.68 56.48 59.81 54.01 54.88 54.32 60.76
ModelScope 52.45 54.48 44.80 45.01 56.70 60.34 53.20 54.42 55.09 58.17
ZeroScope-576w 51.07 54.09 43.36 43.82 55.98 58.74 54.55 54.68 50.39 59.12
LVDM 45.80 46.04 44.45 44.64 40.44 43.09 53.68 53.26 44.61 43.16
Table 8: Performace of T2V models on EvalCrafter with and without best-of-5 sampling using VideoScore. Most EvalCrafter scores have increased compared to the random sample, proving the effectiveness of VideoScore
@@ 4.6 Ablation Study
To investigate the effects of changing the base model, we choose VideoLLaVA-7B and Idefics2-8B as base models from recent popular VLMs (Lin et al. (2023), Laurençon et al. (2024), Zhang et al. (2023), Li et al. (2024a)). Since VideoFeedback-test, EvalCrafter, and VBench both have multiple aspects in the benchmarks, we take the average score across these aspects and report the general performance in  Table 7. The results show that the Video-LLaVA-based version gets the worst performance on the four benchmarks, even if it is specifically designed for video understanding. The Idefics2-8B-based version has marginal improvements compared to the VideoLLaVA. After changing to Mantis-Idefics2-8B, the scores on the four benchmarks keep improving from $47.5$ to $55.6$ on average. When the scoring type is regression, the mantis-based version is still better than the Idefics2-based version by $12.1$ points. Therefore, we select the Mantis-based version as the final choice.
@@ Appendix F Annotation Details
Visual Quality Expected Case: (1) The video looks clear and normal on its appearance. (2) The features like Brightness, Contrast, Color, etc, are appropriate and stable. Error point: (a) local obvious unclear or blurry, (b) too low resolution, (c) some speckles or black patches, (d) appearance of video is skewed and distorted, (e) unstable optical property, such as brightness, contrast, saturation, exposure etc, (f) flickering color of main objects and background Note:Some videos have watermark, we can ignore that.
Temporal Consistency Expected Case: (1) The main objects, main characters and overall appearance are consistent across the video. (2) The appearance of video as well as the movements of humans and objects are smooth and natural. Error points: (a) The person or object suddenly disappears or appears, (b) The type or class of objects has obvious changes, (c) There is an obvious switch in the screen shot,(d) the appearance of video or movements in it is laggy and un-smooth, (e) local deformation or dislocation of human or objects due to the motion. (for large scale deformation, the video should also be rated as bad in "1. visual quality"). Note: For a video almost static or with small dynamic degree, as long as it does not have error points, then it should be scored as good.
Dynamic Degree Expected Case: (1) The video is obviously not static, the people or objects or the video screen is dynamic. (2) The video can be easily distinguished from a static image. Note: You are supposed to focus on only dynamic degree, regardless of the visual quality and video content
Text-to-Video Alignment Expected Case: The characters, objects, motions, events etc. that are mentioned in text input prompts all exist reasonably. Error points: (a) The people and objects in prompt do not appear in video, (b) The actions and events in prompt do not appear in video, (c) The number, size, shape, color, state, movement and other attributes of the objects in the video do not match the prompt, (d) Text mentioned in prompt is not displayed correctly in the video, such as "a placard saying ’No Smoking’" but "No Smoking" is not spelled correctly in the video, (e) The video format (such as width, height, screen ratio, duration) does not match the format in prompt.
Factual Consistency Expected Case: (1) Overall appreance and motion are consistent with our common-sense, physical principles, moral standards, etc. Error points: (a) static ones: Content in video goes against common sense in life, such as lighting a torch in the water, standing in the rain but not getting wet, etc. (b) static ones: The size, color, shape and other basic properties of objects violate scientific principles (c) dynamic ones: The overall movement of people or objects violates common-sense and laws of physics, such as spontaneous upward movement against gravity, abnormal water flow, etc. (d) dynamic ones: Partial movements of people or objects violate common-sense and laws of physics, such as the movement of hands or legs is anti-joint, etc. Notes: Relation with ’5. text-to-video alignment’: Some text prompts express fictional and unrealistic content, for example, "a dog plays the guitar in the sky" or "an astronaut rides a horse in space". In this case, regardless of the veracity of the text prompt, you should only consider whether the other content in the video makes sense.
@@ Appendix I Case study of VideoFeedback
Appendix J Leaderboard
We generated 200 videos using various T2V models with prompts sampled from our prompt set, then we rank these T2V models based on the average score of five dimensions in VideoScore, as shown in  Table 13.
T2V Model Avg VQ TC DD TVA FC
Kling 81.83 84.63 82.49 85.10 74.29 82.65
OpenSora-v1.2 69.85 70.36 67.17 75.91 68.17 67.62
Morph Studio 67.56 66.79 70.98 66.71 68.55 64.77
Morph Studio 67.56 66.79 70.98 66.71 68.55 64.77
VideoCrafter-2 66.23 66.72 68.81 67.04 65.77 62.79
Gen-2 65.74 65.14 67.37 67.97 65.39 62.84
PikaLab 65.72 67.21 67.98 65.17 63.79 64.43
Video-LaVIT 64.04 67.57 59.68 70.83 66.7 55.40
LaVie-base 62.86 63.11 59.33 70.25 62.40 59.20
MagicTime 62.61 65.05 63.2 67.33 61.99 55.50
HotShot-XL 62.31 58.55 59.63 70.90 63.93 58.54
Latte 62.02 63.34 59.23 68.38 62.56 56.60
OpenSora-v1.0 61.94 62.49 58.75 69.65 64.30 54.51
OpenSora-v1.1 61.29 61.50 56.82 69.82 64.72 53.61
VideoCrafter-1-512 60.92 60.94 60.12 67.38 60.44 55.73
AnimateDiff 57.33 64.83 41.89 73.46 65.76 40.69
ModelScope 52.45 47.41 49.76 69.14 54 41.94
LVDM 47.20 33.15 44.01 72.75 58.85 27.25
ZeroScope_576w 44.69 31.35 39.65 73.76 49.40 29.27
Table 13: Leaderboard of VideoScore for existing text-to-video models on 200 curated examples.